OpenStudy (anonymous):

If 3x^2-2x+7=0 than what does (x-1/3)^2 equal?

OpenStudy (anonymous):

Solved this one too (x-1/3)^2=x^2-2/3x+1/9 Divide the original equation by 3 and you get x^2-2x+7/3=0 If you add 7/3 and get zero, then if you only add 1/9 you'd get -20/9